OpenStudy (anonymous):

x+y=10 -x+y=0 (answer by substitution)

OpenStudy (alexandervonhumboldt2):

x=10-y x=y

OpenStudy (alexandervonhumboldt2):

y=10-y

OpenStudy (alexandervonhumboldt2):

2y=10 y=5

OpenStudy (alexandervonhumboldt2):

x=5

OpenStudy (alexandervonhumboldt2):

(5, 5)

OpenStudy (anonymous):

@AlexandervonHumboldt2 dammmm you are fast! thanks though